# Unlocking Multi-Dimensional Integration with Quantum Adaptive Importance Sampling

### v1

@@ 4.2 Multi-peak Benchmark Integrals
where $\mathbf{r_{0}}=(0.23,\dots,0.23)$, $\mathbf{r_{1}}=(0.39,\dots,0.39)$, and $\mathbf{r_{2}}=(0.74,\dots,0.74)$. We consider the cases $d=2,3,4$, using the same discretization in all instances, which is five qubits per dimension. For the optimization, we employed $20\times 10^{3}$ to $85\times 10^{3}$ iterations for the PQC with $225$-$750$ parameters. The Ansatz chosen is the sequence $U^{(Z)}\to U_{3}\to U^{(Y)}\to U_{3}\to U^{(X)}\to U_{3}$, which proved to be the best choice while considering the final KL divergence and the number of iterations. The resulting KL divergences fall in the range $0.09-0.21$.
For generating the results, we made multiple identical integration runs and report the average uncertainty. It shall be noted that we observed a rare effect in the four-dimensional case, which we attribute to insufficient coverage of the Important Region and the imperfect training. In these rare runs, the variance would be getting considerably inflated. This was a consequence of not adequate sampling of one of the peaks in the extended boundary region. Thus, a large tile would be generated and evaluated with a single sample, which could fall to a badly covered part in proximity to the Important Region. To counteract this effect, we use a defensive IS mixture, to stabilize such a behavior, a strategy that is considered unbiased hesterberg1995weighted . The mixture we have used is to sample $90\%$ of the points from the quantum-generated proposal PDF, and the remaining $10\%$ from a uniform PDF that are distributed in the grid after the QAIS proposal has generated it.
@@ 5 Conclusions and Outlook
For the demonstration of the performance of the method, we applied QAIS to integrate two challenging types of functions and compared the results with VEGAS, which serves as our reference. These functions are a sharply-peaked single-mode one-loop Feynman integrand, and a set of multi-modal benchmark integrands from two to four dimensions. For the first case, the two methods performed comparably overall. For the multi-modal functions, VEGAS allocated a substantial amount of samples in non-Important Regions, especially as the dimension increased, due to its separable proposal PDF, whereas QAIS consistently maintained impressive accuracy.
@@ Appendix A Tiling Algorithm
where each $p(s_{k})$ is the empirical frequency of $s_{k}$. Sorting these strings in an ascending order costs $\mathcal{O}(M\log M)$, or $\mathcal{O}(n~{}M\log M)$ if the comparison cost in $n$ bits is considered also. For consecutive sorted strings $s_{k}<s_{k+1}$ we define the gap size :
@@
If $\Delta_{k}>0$ there are $\Delta_{k}$ unobserved cells strictly between $s_{k}$ and $s_{k+1}$. If $\Delta_{k}=0$ the two observed cells are adjacent. Then, we proceed with the greedy construction of the hyper-rectangles to cover the missing ranges. The main difficulty is to cover every gap with a few hyper-rectangles, depending only on $d$ and $n$ and the number of measured outcomes $M$ while not depending on the exponentially large pool of unmeasured states $2^{n}-M=\sum_{k}\Delta_{k}$. The main procedure (lines 26-44 of Algorithm 1) expands greedily from the least-significant right-most dimension towards the most-significant left-most one, grouping whole blocks of cells whenever lower-order dimensions are already fully spanned.
Starting from the lower corner of a gap, the tiling algorithm iteratively constructs a tile that never exceeds the boundary $2^{q_{i}}$ on any axis, never exceeds the remaining gap size $\Delta_{k}$, and expands the dimension $i+1$ only after dimension $i$ has been filled from coordinate 00 up to its boundary. In each dimension except the highest-order one, at most two separate expansions are made. The top dimension cannot have full expansions. Consequently, a single gap is covered by at most $12(d-1)+$ 1 hyper-rectangles.
Once all gaps are tiled, we obtain an ordered list of measured points and hyper-rectangles that jointly cover the entire space. Collapsing the unmeasured tiles into the non-Important Region (as described in Sec. 3.3) is now straightforward. Each empty tile is merged with the next measured state, the latter’s probability $p(s_{k+1})$ being assigned to the entire hyper-rectangle. If necessary, for the last interval, ( from $s_{M}$ to the binary state $11\dots 1$ 11 ) , we generate an additional tile, and merge it into the group containing $s_{M}$. Because the associated list is compact and sorted, alternative ways of performing the de-biasing are possible. For example, merging neighboring tiles with common boundaries.
@@
1:Input:     $\bullet$ Grid dimensions: $grid\_dims=[2^{q_{d}},\ldots,2^{q_{1}}]$
2:                $\bullet$ Dictionary of Measured States as Bitstrings and Probabilities: $S$
3:Output: Dictionary, with tiles (in coordinate format), that cover the integration space.
4:==================================
5:Initialize Result Dictionary (Associated Array): $F\leftarrow\{\}$
6:Sort the states: $S\leftarrow\left\{(s_{1},p(s_{1})),(s_{2},p(s_{2})),\dots,(s_{M},p(s_{M}))\right\}$triangleright $\mathcal{O}(M\log M)$
7:for $i\leftarrow 1$ to $M-1$ $\triangleright$ $\mathcal{O}(M)$
8:    $\Delta\leftarrow\text{int}(s_{i+1})-\text{int}(s_{i})-1$triangleright Cells to fill
9:    if $\Delta>0$ then
10:        $current\_linear\leftarrow\text{int}(s_{i})+1$
11:        $tile\_intervals\leftarrow\textsc{IntervalsGenerator}(grid\_dims,current\_linear,\Delta)$triangleright $\mathcal{O}(d^{3}n^{2})$
12:        Append $(tile\_intervals,0)$ to $F$
13:    else if $\Delta=0$ then
14:        $end\_coord\leftarrow current\_coord$
15:        Append $((current\_coord,end\_coord),p(s_{i}))$ to $F$
@@
17:end for
18:return $F$
19:==================================
20:procedure IntervalsGenerator($grid\_dims,current\_linear,\Delta$)
21:    Initialize: $intervals\leftarrow\emptyset$
22:    while $\Delta>0$ do $\triangleright$ # of loops bounded by: $|intervals|\approx\mathcal{O}(d)$
23:        $current\_coord\leftarrow lin\_to\_coord(current\_linear)$triangleright Int. to grid coord. (BEE)
24:        $(tile\_shape,cells\_filled)\leftarrow\textsc{GreedyMultidimExpand}(current\_coord,\ \Delta)$
25:        $end\_coord\leftarrow\left(current\_coord[i]+tile\_shape[i]-1\right),\quad i=1,\dots,d$
26:        Append $(current\_coord,end\_coord)$ to $intervals$
27:        $\Delta\leftarrow\Delta-cells\_filled$
28:        $current\_linear\leftarrow current\_linear+cells\_filled$
29:    end while
30:    return $intervals$triangleright $|intervals|\leq 2*(d-1)+1$
31:end procedure
32:==================================
33:procedure GreedyMultidimExpand($current\_coord,\Delta$)
34:    Initialize: $tile\_shape\leftarrow[1,\ldots,1]$triangleright $|tile\_shape|=d$
35:    $tile\_shape[d]\leftarrow\min\{\Delta,grid\_dims[d]-current\_coord[d]\}$
36:    for $i\leftarrow d-1$ down to $1$ do $\triangleright$ $\mathcal{O}(d)$
37:        if $current\_coord[i+1]=0$ and $tile\_shape[i+1]=grid\_dims[i+1]$ then
38:           $comb\_vol\leftarrow\prod_{j=i+1}^{d}grid\_dims[j]$triangleright $\mathcal{O}(d\,n^{2})$
39:           $tile\_extend\leftarrow\lfloor\Delta/comb\_vol\rfloor$
40:           if $tile\_extend>0$ then
41:               $tile\_shape[i]\leftarrow\min\{grid\_dims[i]-current\_coord[i],tile\_extend\}$
42:           else
43:               $tile\_shape[i]\leftarrow 1$
44:           end if
45:        else
46:           $tile\_shape[i]\leftarrow 1$
47:        end if
48:    end for
49:    $block\_vol\leftarrow\prod_{i=1}^{d}tile\_shape[i]$
50:    return $(tile\_shape,block\_vol)$
51:end procedure
52:==================================

### v2

@@ 4.2 Multi-peak Benchmark Integrals
where $\mathbf{r_{0}}=(0.23,\dots,0.23)$, $\mathbf{r_{1}}=(0.39,\dots,0.39)$, and $\mathbf{r_{2}}=(0.74,\dots,0.74)$. We consider the cases $d=2,3,4$, using the same discretization in all instances, which is five qubits per dimension. For the optimization, we employed $20\times 10^{3}$ to $85\times 10^{3}$ iterations for the PQC with 225225225-750750750 parameters. The Ansatz chosen is the sequence $U^{(Z)}\to U_{3}\to U^{(Y)}\to U_{3}\to U^{(X)}\to U_{3}$, which proved to be the best choice while considering the final KL divergence and the number of iterations. The resulting KL divergences fall in the range 0.090.210.09-0.210.09 - 0.21.
[…]
@@ 5 Conclusions and Outlook
For the demonstration of the performance of the method, we applied QAIS to integrate two challenging types of functions and compared the results with VEGAS, which serves as our reference. The comparison focuses on the accuracy that each method’s optimal proposal PDF can achieve. The functions considered are a sharply-peaked single-mode one-loop Feynman integrand, and a set of multi-modal benchmark integrands from two to four dimensions. For the first case, the two methods performed comparably overall. For the multi-modal functions, VEGAS allocated a substantial amount of samples in non-Important Regions, especially as the dimension increased, due to its separable proposal PDF, whereas QAIS consistently maintained impressive accuracy.
@@ Appendix A Tiling Algorithm
where each $p(s_{k})$ is the empirical frequency of $s_{k}$. If the first and last possible states (namely the binary states $\dots 00$ 00 and $\dots 11$ 11) have not been observed, we manually insert them with zero probability. Sorting these strings in an ascending order costs $\mathcal{O}(M\log M)$, or $\mathcal{O}(n~M\log M)$ if the comparison cost in $n$ bits is considered also. For consecutive sorted strings $s_{k}<s_{k+1}$ we define the gap size:
@@
If $\Delta_{k}>0$ there are $\Delta_{k}$ unobserved cells strictly between $s_{k}$ and $s_{k+1}$. If $\Delta_{k}=0$ the two observed cells are adjacent. Then, we proceed with the greedy construction of the hyper-rectangles to cover the missing ranges. The main difficulty is to cover every gap with a few hyper-rectangles, depending only on $d$ and $n$ and the number of measured outcomes $M$ while not depending on the exponentially large pool of unmeasured states $2^{n}-M=\sum_{k}\Delta_{k}$. The main procedure (lines 32-50 of Algorithm 1) expands greedily from the least-significant right-most dimension towards the most-significant left-most one, grouping whole blocks of cells whenever lower-order dimensions are already fully spanned.
Starting from the lower corner of a gap, the tiling algorithm iteratively constructs a tile that never exceeds the boundary $2^{q_{i}}$ on any axis, never exceeds the remaining gap size $\Delta_{k}$, and expands the dimension $i+1$ only after dimension $i$ has been filled from coordinate 0 up to its boundary. In each dimension except the highest-order one, at most two separate expansions are made. The top dimension cannot have full expansions. Consequently, a single gap is covered by at most $2(d-1)+1$ hyper-rectangles.
Once all gaps are tiled, we obtain an ordered list of measured points and hyper-rectangles that jointly cover the entire space. Collapsing the unmeasured tiles into the non-Important Region (as described in Sec. 3.3) is now straightforward. Each empty tile is merged with the next measured state, the latter’s probability $p(s_{k+1})$ being assigned to the entire hyper-rectangle. If necessary, for the last interval, ( from $s_{M-1}$ to the binary state $\dots 11$ 11 ), we merge it into the group containing $s_{M-1}$. This is the case when the final state is unobserved. Because the associated list is compact and sorted, alternative ways of performing the de-biasing are possible. For example, merging neighboring tiles with common boundaries.
@@
1:Input:
  • Grid dimensions: $grid\_dims\leftarrow[2^{q_{d}},\ldots,2^{q_{1}}]$
  • Dictionary of Measured States as Bitstrings and Probabilities : $S$
2:Output: Dictionary, with tiles (in coordinate format), that cover the integration space.
3:=========================================================
4:Initialize Result Dictionary (Associated Array): ${}F\leftarrow\{\}$
5:Sort the states: $S\leftarrow\left\{(s_{1},p(s_{1})),(s_{2},p(s_{2})),\dots,(s_{M},p(s_{M}))\right\}$triangleright $\mathcal{O}(n\;M\log M)$
6:for $i\leftarrow 1$ to $M$ $\triangleright$ $\mathcal{O}(M)$
7:  $current\_coord\leftarrow lin\_to\_coord(\text{int}(s_{i}))$triangleright To integer, to grid coordinate (BEE)
8:  $end\_coord\leftarrow current\_coord$
9:  Append $((current\_coord,end\_coord),p(s_{i}))$ to $F$
10:  if $i<M$ then
11:   $\Delta\leftarrow\text{int}(s_{i+1})-\text{int}(s_{i})-1$triangleright Cells to fill
12:   if $\Delta>0$ then
13:     $current\_linear\leftarrow\text{int}(s_{i})+1$
14:     $tile\_intervals\leftarrow\textsc{IntervalsGenerator}(grid\_dims,current\_linear,\Delta)$triangleright $\mathcal{O}(d^{3}n^{2})$
15:     Append $(tile\_intervals,0)$ to $F$
@@
17:  end if
18:end for
19:return $F$
20:=========================================================
[…]
22:  Initialize: $intervals\leftarrow\emptyset$
23:  while $\Delta>0$ do $\triangleright$ # of loops bounded by: $|intervals|\approx\mathcal{O}(d)$
24:   $current\_coord\leftarrow lin\_to\_coord(current\_linear)$
25:   $(tile\_shape,cells\_filled)\leftarrow\textsc{GreedyMultidimExpand}(current\_coord,\ \Delta)$triangleright $\mathcal{O}(d^{2}\,n^{2})$
26:   $end\_coord\leftarrow\left(current\_coord[i]+tile\_shape[i]-1\right),\quad i=1,\dots,d$
27:   Append $(current\_coord,end\_coord)$ to $intervals$
28:   $\Delta\leftarrow\Delta-cells\_filled$
29:   $current\_linear\leftarrow current\_linear+cells\_filled$
30:  end while
31:  return $intervals$triangleright $|intervals|\leq 2*(d-1)+1$
32:end procedure
33:=====================================================
[…]
35:  Initialize: $tile\_shape\leftarrow[1,\ldots,1]$triangleright $|tile\_shape|=d$
36:  $tile\_shape[d]\leftarrow\min\{\Delta,grid\_dims[d]-current\_coord[d]\}$
37:  for $i\leftarrow d-1$ down to 111 do $\triangleright$ $\mathcal{O}(d)$
38:   if $current\_coord[i+1]=0$ and $tile\_shape[i+1]=grid\_dims[i+1]$ then
39:     $comb\_vol\leftarrow\prod_{j=i+1}^{d}grid\_dims[j]$triangleright $\mathcal{O}(d\,n^{2})$
40:     $tile\_extend\leftarrow\lfloor\Delta/comb\_vol\rfloor$
41:     if $tile\_extend>0$ then
42:      $tile\_shape[i]\leftarrow\min\{grid\_dims[i]-current\_coord[i],tile\_extend\}$
43:     else
44:      $tile\_shape[i]\leftarrow 1$
45:     end if
46:   else
47:     $tile\_shape[i]\leftarrow 1$
48:   end if
49:  end for
50:  $block\_vol\leftarrow\prod_{i=1}^{d}tile\_shape[i]$
51:  return $(tile\_shape,block\_vol)$
52:end procedure
53:=====================================================